OpenStudy (anonymous):

3x-y=5 7x-2y=13 what is the solution?

OpenStudy (compassionate):

We guide you to out answers here at OpenStudy. We guide you to the solution. What do you think the first step is?

OpenStudy (anonymous):

do i multiply 3x5?

OpenStudy (compassionate):

It depends. Is this a system of equations or two separate equations?

OpenStudy (anonymous):

system of equations

OpenStudy (compassionate):

3x-y=5 7x-2y=13 Start off by solving the first equation for y. Can you do that?

OpenStudy (anonymous):

iam not very good at math, iam good at the basics like multiply,divide,adding,subtracting

OpenStudy (compassionate):

Alright. So solve an equation you want to do the oppsite. 3x - y = 5 So, we subtract 3x -y = -3x + 5 Now, y cannot be negative so we divide y = 3x - 5 Got it? Now, we know y = 3x - 5, so plug that into the equation below it. 7x-2(3x - 5) =13 <-- now solve for x.

OpenStudy (compassionate):

http://www.mathway.com/ solves equations for you.

OpenStudy (anonymous):

1. chose any of the two equations that your are comfortable with in finding "x" 2. in this situation, i chose 3x-y=5 3x - y = 5 3x = 5 +y (transpose y to the other side) x = 5/3 + y/3 (divide both sides by 3) 3. now, substitute the value of "x" to the second equation 7x - 2y = 13 7(5/3 + y/3) - 2y = 13 4. multiply the one inside in the parenthesis by 7: 35/3 + 7y/3 - 2y = 13 7y/3 - 2y = 13 - 35/3 (transpose 35/3 to the other side) 5. since the common denominator is 3, multiply both sides by 3 now it will become; 7y -6y = 39 -35 y = 4 6. substitute to either of the two equations: i chose 3x - y = 5 3x - (4) = 5 (substitute "y" to "4") 3x = 5 +4 (transpose 4 to the other side) 3x = 9 x = 3 (divide both sides by 3) the answer is x=3 and y=4 how to know if your solution is correct? substitute either of the two equations the values of "x" and "y" i chose 7x - 2y = 13 7(3) - 2(4) = 13 21 - 8 = 13 13 = 13 therefore, the solution is correct....